# LinearRAG: Linear Graph Retrieval Augmented Generation on Large-scale Corpora

### v1

@@ Appendix B Baseline Details
(i) KGP (Wang et al., 2024) builds a knowledge graph over multiple passages with LLMs; at retrieval stage, it introduces an LLM-driven graph traversal agent to navigate the graph and progressively collect supporting passage.
(ii) G-Retriever (He et al., 2024) combines graph neural networks with LLMs by formulating subgraph retrieval as a Prize-Collecting Steiner Tree optimization problem, enabling effective conversational question answering on textual graphs while mitigating hallucination and enhancing scalability.
(iii) RAPTOR (Sarthi et al., 2024) develops a hierarchical tree by applying clustering algorithms and abstractive summarization techniques, facilitating representation at multiple semantic granularities.
(iv) E2GraphRAG (Zhao et al., 2025) uses spaCy to extract entities and LLMs to summarize passage groups into a hierarchical tree with encoded nodes at indexing stage; at retrieval stage, it hits relevant entities in $k$-hop neighborhood and collects associated passages for ranking, otherwise dense retrieval over the whole tree.
(v) LightRAG (Guo et al., 2024) employs a two-tier framework that incorporates graph-based representations within textual indexing, merging fine-grained entity-relation mappings with coarse-grained thematic structures.
(vi) HippoRAG (Gutiérrez et al., 2024) is a training-free graph-enhanced retriever that uses the Personalized PageRank algorithm with query concepts as seeds for single-step or multi-hop retrieval across disparate documents.
(vii) GFM-RAG (Luo et al., 2025) implements a GraphRAG paradigm by constructing graphs from documents and using a graph-enhanced retriever to retrieve relevant documents.
(viii) HippoRAG2 (Gutiérrez et al., 2025) extends HippoRAG with enhanced paragraph integration and contextualization. Optimizes seed node selection and PageRank reset probabilities while maintaining factual memory capabilities and improving associative memory performance. improving associative memory performance.
@@ E.3 Effectiveness of Different Sentence Embeddings (Q6)
In this section, we study the effectiveness of different sentence embeddings in the LinearRAG. We compare the all-mpnet-base-v2 (Song et al., 2020), bge-large-en (Xiao et al., 2023), all-MiniLM-L6-v2 (Wang et al., 2020) and e5-large-v2 (Wang et al., 2022). We download the official pre-trained model from the Huggingface2.
@@ E.4 Large-Scale Efficiency Analysis (Q7)
To evaluate the scalability and effectiveness of LinearRAG on large corpora, we conduct comprehensive efficiency analysis using the ATLAS-Wiki dataset (Bai et al., 2025). We create two subsets containing 5M and 10M tokens respectively for indexing phase evaluation, as shown in Table 6.
@@ Appendix F Related Work
Large language models are prone to hallucination (Fang et al., 2024; Jiang et al., 2025; Fang et al., 2025; Zheng et al., 2025; Hong et al., 2024; Yuan et al., 2025; Zhong et al., 2024), while RAG is a promising solution by grounding the reasoning process on contextual evidence from knowledge bases (Zhang et al., 2025a; Zhou et al., 2025b; Zhang et al., 2025b; Xiang et al., 2025). However, real-world knowledge is often distributed across documents, organizing them effectively for answering complex questions has always been a challenging yet promising research topic in RAG. In the following, we discuss two major lines (❶, ❷) of graphRAG research that closely relate to our work, which construct explicit graphs for organizing external knowledge sources. Then discuss reasoning-enhanced RAG (❸), which leverages LLM’s inherent reasoning ability for answering complex questions without explicitly structuring corpora.
@@
Clustering-based hierarchy construction. One line of research employs clustering-based community detection, a bottom-up method that applies algorithms like Louvain or Leiden to identify densely connected clusters of entities within the initial graph (Edge et al., 2024; Sarthi et al., 2024; Gutiérrez et al., 2024). By grouping related entities, it creates a hierarchical structure that abstracts passages into topic-based communities, reducing redundancy and offering a broader, macro-level view of the data. However, as an unsupervised technique, it is vulnerable to error propagation, where inaccuracies in entity relationships are amplified at higher levels of abstraction. Moreover, applying these clustering algorithms to large-scale graphs presents significant scalability issues, making real-time applications infeasible in practice.
@@
Relation-extraction-based knowledge graph construction. This line of research (Gutiérrez et al., 2024; Guo et al., 2024; Xiao et al., 2025c) adopts the idea of knowledge graphs to organize knowledge across different passages. The basic idea is to extract triples from each text chunk (passage) as an atomic summarization of the knowledge within the passage. The triples are connected via entity alignment (Chen et al., 2024a; b), ultimately forming a unified knowledge graph, which directly carries structured knowledge and serves as the index of passages, where off-the-shelf graph reasoning algorithms (Shengyuan et al., 2024; Qu & Tang, 2019; Liu et al., 2023) can be applied to perform reasoning (Sun et al., 2024; LUO et al., 2024). However, due to context window limits, OpenIE on each passage is processed independently, so the generated triples may be inconsistent. Recent work (Liang et al., 2024; Sharma et al., 2024) mitigates this via top-down graph construction guided by a globally defined schema, but it relies on manual expert annotation to create and maintain the domain-specific schema, which is not only expensive, time-consuming, and does not generalize well across domains.
@@
Reasoning-enhanced RAG. The key to answering complex questions in RAG is to effectively identify multiple distributed documents that support multihop logical dependencies. While GraphRAG methods explicitly construct graphs with human priors to organize corpora for efficient structured retrieval, reasoning-enhanced RAG directly leverages the inherent reasoning ability of LLMs to decompose complex queries into simpler subquestions that can be addressed by dense retrieval. LogicRAG (Chen et al., 2025) and LAG (Xiao et al., 2025b) formalize this idea by first decomposing a query into a set of subqueries, then constructing a directed acyclic graph based on the logical dependencies among them, and solving subproblems one by one by following the topological order. Similarly, Chain-of-Note (CoN) (Yu et al., 2024) enhances RAG by generating sequential notes that break down complex queries into intermediate reasoning steps, retrieving relevant documents for each step to build a coherent answer. Another approach, SelfRAG (Asai et al., 2024), integrates self-reflection into the RAG process, where the LLM iteratively evaluates and refines subqueries to ensure retrieved documents align with the logical flow of the reasoning process. These methods collectively highlight the potential of LLM-driven reasoning to improve retrieval efficiency and answer quality for complex, multihop queries.

### v4

@@ Appendix B Baseline Details
(i) KGP (wang2024knowledge) builds a knowledge graph over multiple passages with LLMs; at retrieval stage, it introduces an LLM-driven graph traversal agent to navigate the graph and progressively collect supporting passage.
(ii) G-Retriever (he2024g) combines graph neural networks with LLMs by formulating subgraph retrieval as a Prize-Collecting Steiner Tree optimization problem, enabling effective conversational question answering on textual graphs while mitigating hallucination and enhancing scalability.
(iii) RAPTOR (sarthi2024raptor) develops a hierarchical tree by applying clustering algorithms and abstractive summarization techniques, facilitating representation at multiple semantic granularities.
(iv) E2GraphRAG (zhao20252GraphRAG) uses spaCy to extract entities and LLMs to summarize passage groups into a hierarchical tree with encoded nodes at indexing stage; at retrieval stage, it hits relevant entities in $k$-hop neighborhood and collects associated passages for ranking, otherwise dense retrieval over the whole tree.
(v) LightRAG (guo2024lightrag) employs a two-tier framework that incorporates graph-based representations within textual indexing, merging fine-grained entity-relation mappings with coarse-grained thematic structures.
(vi) HippoRAG (hipporag) is a training-free graph-enhanced retriever that uses the Personalized PageRank algorithm with query concepts as seeds for single-step or multi-hop retrieval across disparate documents.
(vii) GFM-RAG (luo2025gfm) implements a GraphRAG paradigm by constructing graphs from documents and using a graph-enhanced retriever to retrieve relevant documents.
(viii) HippoRAG2 (gutiérrez2025hipporag2) extends HippoRAG with enhanced paragraph integration and contextualization. Optimizes seed node selection and PageRank reset probabilities while maintaining factual memory capabilities and improving associative memory performance. improving associative memory performance.
@@ E.3 Effectiveness of Different Sentence Embeddings (Q6)
In this section, we study the effectiveness of different sentence embeddings in the LinearRAG. We compare the all-mpnet-base-v2 (song2020mpnet), bge-large-en (bge_embedding), all-MiniLM-L6-v2 (wang2020minilm) and e5-large-v2 (wang2022text). We download the official pre-trained model from the Huggingface2.
@@ E.4 Large-Scale Efficiency Analysis (Q7)
To evaluate the scalability and effectiveness of LinearRAG on large corpora, we conduct comprehensive efficiency analysis using the ATLAS-Wiki dataset (bai2025autoschemakg). We create two subsets containing 5M and 10M tokens respectively for indexing phase evaluation, as shown in Table 6.
@@ Appendix F Related Work
Large language models are prone to hallucination (fang2024alphaedit; jiang2025anyedit; fang2025safemlrm; zheng2025usb; hong2024next; yuan2025knapsack; zhong2024iterative), while RAG is a promising solution by grounding the reasoning process on contextual evidence from knowledge bases (zhang2025erarag; zhou2025depth; xiao2025graphragbenchchallengingdomainspecificreasoning; zhang2025survey; xiang2025use). However, real-world knowledge is often distributed across documents, organizing them effectively for answering complex questions has always been a challenging yet promising research topic in RAG. In the following, we discuss two major lines (❶, ❷) of graphRAG research that closely relate to our work, which construct explicit graphs for organizing external knowledge sources. Then discuss reasoning-enhanced RAG (❸), which leverages LLM’s inherent reasoning ability for answering complex questions without explicitly structuring corpora.
@@
Clustering-based hierarchy construction. One line of research employs clustering-based community detection, a bottom-up method that applies algorithms like Louvain or Leiden to identify densely connected clusters of entities within the initial graph (edge2024local; sarthi2024raptor; hipporag). By grouping related entities, it creates a hierarchical structure that abstracts passages into topic-based communities, reducing redundancy and offering a broader, macro-level view of the data. However, as an unsupervised technique, it is vulnerable to error propagation, where inaccuracies in entity relationships are amplified at higher levels of abstraction. Moreover, applying these clustering algorithms to large-scale graphs presents significant scalability issues, making real-time applications infeasible in practice.
@@
Relation-extraction-based knowledge graph construction. This line of research (hipporag; guo2024lightrag; xiao2025reliablereasoningpathdistilling) adopts the idea of knowledge graphs to organize knowledge across different passages. The basic idea is to extract triples from each text chunk (passage) as an atomic summarization of the knowledge within the passage. The triples are connected via entity alignment (neusymea; LLM4EA), ultimately forming a unified knowledge graph, which directly carries structured knowledge and serves as the index of passages, where off-the-shelf graph reasoning algorithms (difflogic; pLogicNet; liu2023rsc) can be applied to perform reasoning (sun2024thinkongraph; luo2024reasoning). However, due to context window limits, OpenIE on each passage is processed independently, so the generated triples may be inconsistent. Recent work (liang2024kag; sharma2024og) mitigates this via top-down graph construction guided by a globally defined schema, but it relies on manual expert annotation to create and maintain the domain-specific schema, which is not only expensive, time-consuming, and does not generalize well across domains.
@@
Reasoning-enhanced RAG. The key to answering complex questions in RAG is to effectively identify multiple distributed documents that support multihop logical dependencies. While GraphRAG methods explicitly construct graphs with human priors to organize corpora for efficient structured retrieval, reasoning-enhanced RAG directly leverages the inherent reasoning ability of LLMs to decompose complex queries into simpler subquestions that can be addressed by dense retrieval. LogicRAG (chen2025logicrag) and LAG (xiao2025lag) formalize this idea by first decomposing a query into a set of subqueries, then constructing a directed acyclic graph based on the logical dependencies among them, and solving subproblems one by one by following the topological order. Similarly, Chain-of-Note (CoN) (chainofNote) enhances RAG by generating sequential notes that break down complex queries into intermediate reasoning steps, retrieving relevant documents for each step to build a coherent answer. Another approach, SelfRAG (asai2024selfRAG), integrates self-reflection into the RAG process, where the LLM iteratively evaluates and refines subqueries to ensure retrieved documents align with the logical flow of the reasoning process. These methods collectively highlight the potential of LLM-driven reasoning to improve retrieval efficiency and answer quality for complex, multihop queries.